### v1

# Impact of Reactive Jamming Attacks on LoRaWAN: a Theoretical and Experimental Study

Abstract
This paper investigates the impact of reactive jamming on LoRaWAN networks, focusing on minimizing jammer exposure time while effectively disrupting communication. We analyze the protection mechanisms implemented in LoRa and explore how different jamming configurations influence frame success rates. A key contribution of this work is the proposal of a Software Defined Radio (SDR)-based jamming approach that generates a controlled number of random symbols, independent of the standard LoRa frame structure. This approach enables precise control over jammer exposure time and provides flexibility in studying the effect of jamming symbols on network performance. Theoretical analysis is validated through experimental results, where a jammer implemented on GNU Radio is used to assess the impact of jamming under various configurations. Our findings demonstrate that LoRa-based networks can be disrupted with a minimal number of symbols, emphasizing the need for future research on covert communication techniques to counter such jamming attacks.
Index Terms:
LoRaWAN, LoRa, jamming, denial-of-service, GNU Radio.
I Introduction
The Internet of Things (IoT) has transformed numerous sectors, including healthcare, agriculture, and smart cities, by enabling seamless communication and data sharing among interconnected devices. At the heart of IoT systems lies the communication protocol, which is essential for data transfer between devices. IoT networks consist of resource-constrained devices that must communicate over long distances. However, most existing communication technologies are either limited in range, such as Zigbee and Bluetooth, or energy-intensive, like 802.11x [1]. This led to the emergence of Low Power Wide Area Network (LPWAN) technologies, including Long Range (LoRa) and its network protocol LoRaWAN (Long Range Wide Area Network), which have gained popularity due to their long-range capabilities and cost-effectiveness.
LoRaWAN’s increasing adoption has introduced several challenges, particularly related to scalability and security [2]. As the number of devices grows, packet collisions rise, reducing network capacity due to the pure Aloha scheme used by LoRaWAN. Various solutions have been proposed to enhance network capacity, such as the adoption of slotted Aloha [3, 4], support for carrier sensing [5], and multi-user receivers with interference cancellation at the gateway [6]. Regarding security, several attacks have been identified, including message replay targeting node activation, packet forging [7, 8], battery exhaustion, and denial-of-service (DoS) attacks [9], along with proposed countermeasures [10].
Among these, DoS attacks from jamming are particularly problematic for the network. They are difficult to fully mitigate and easy to execute, especially since LoRa operates in the ISM (Industrial, Scientific, and Medical) band. However, their impact remains underexplored in the literature. Previous studies on LoRa jamming typically used static setups, where the jammer was configured with a fixed frame length and power solely to demonstrate the effectiveness of jamming.
In this paper, we investigate the impact of jamming on LoRaWAN networks under various configurations. We demonstrate that LoRa-based networks can be jammed with a minimal number of symbols, effectively blocking their traffic, and identify the exact number of symbols required for this disruption. Given that jamming is facilitated by the easily detectable LoRa preamble, we aim for this study to highlight the need for further research into covert communication techniques for LoRa, especially for critical nodes. Our contributions are as follows:
  • We analyze how LoRa’s protection mechanisms behave under jamming and derive expressions to quantify the impact of jamming on frame success rates. Specifically, the derived formulas describe how the number of jamming symbols affects the frame success rate.
  • We propose a Software Defined Radio (SDR)-based jamming approach that generates a controlled number of random symbols, independent of the standard LoRa frame structure. This method enables precise control over jammer exposure time and facilitates experiments on how varying the number of jamming symbols impacts frame success rates.
  • We validate our theoretical analysis through experiments using a jammer implemented on GNU Radio, demonstrating the effects of jamming on LoRa transmissions for various parameters.
The rest of the paper is organized as follows: Section II reviews related work on LoRa jamming. Section III provides an overview of LoRaWAN and LoRa. Section IV discusses LoRa’s protection mechanisms and introduces our jamming framework. Section V presents the experimental validation of our analysis, and Section VI concludes the paper and suggests future work.
II Related Work
The increasing adoption of LoRaWAN has triggered growing concerns about its security, prompting a series of assessments across the different layers of its communication stack. Among these, physical layer jamming attacks have received particular attention due to their ease of implementation and difficulty in mitigation. Such attacks pose a significant threat, potentially resulting in partial or complete denial of service (DoS). Additionally, the operation of LoRa/LoRaWAN on ISM bands has made it a more attractive target for various experimental attacks. As a result, several studies have focused on this type of attack.
In [9], the authors demonstrated an energy depletion attack on a LoRa node, where they first jammed the uplink transmission and then mimicked the gateway by sending a downlink transmission in RX2 to keep the device active. [11] explores the impact of DoS attacks on the energy consumption of LoRaWAN end-devices. [12] provides a detailed analysis of two jamming variants: reactive and selective jamming. In reactive jamming, the attacker listens to the channel and only jams when a frame is detected, whereas selective jamming targets specific devices, leaving others unaffected. [13] investigates the performance of LoRaWAN networks under jamming conditions, evaluating scenarios with and without node retransmission. The study used a network simulator, and the jammer in this case did not listen to the channel, instead randomly transmitting jamming frames.
Several methods have been proposed to detect and/or mitigate jamming attacks on LoRa/LoRaWAN, such as exploiting the differences between jamming signals and legitimate transmissions in the power domain to distinguish between the two [14].
Most of these studies follow a similar experimental setup: a jammer based on a conventional LoRa chip transmits a complete LoRa frame to interfere with ongoing transmissions, increasing the air-time of the jamming signal and the likelihood of detection. However, the effects of jamming frame length and its impact on the inherent LoRa protection mechanisms have not been thoroughly investigated in the literature. This gap is the focus of the present paper.
III Overview of LoRaWAN and LoRa
LoRaWAN is an open-source Low Power Wide Area Network (LPWAN) technology developed and maintained by the LoRa Alliance. It uses the LoRa modulation at the physical layer, complemented by the MAC and network layers to form a complete networking stack. The development and deployment of LoRaWAN adhere to the LoRaWAN specification, which provides a set of rules and recommendations.
The specification defines a star topology, where end devices (ED) communicate with gateways (GW) via radio frequency. The gateways forward uplink frames from the EDs to the network server (NS) and downlink frames from the NS to the EDs. It also outlines three device classes—A, B, and C—each with different power consumption policies and latency characteristics. Class A devices are the most power-efficient, as they remain in sleep mode and only wake up to transmit, waiting to receive during two short windows defined by the specification before returning to sleep. Class B devices can open additional reception windows, while Class C devices are always in reception mode, except when transmitting. Class A is the default mode, required for all LoRaWAN nodes, though it comes with the longest and most unpredictable latency.
The specification also includes physical layer policies, some of which vary by country, such as channel frequency, bandwidth, packet size, power levels, and duty cycle limitations.
LoRa, a proprietary technology developed by Semtech, is designed for low power and long-range communication, making it ideal for Internet of Things (IoT) applications. It uses Chirp Spread Spectrum (CSS) modulation to achieve wide coverage with low data rates. The technology defines six spreading factors (SF) ranging from 7 to 12. The SF, along with bandwidth, determines the air-time of frames: smaller SF values reduce air-time and range, lower power consumption, and increase data rates; larger SF values increase air-time and range, but also raise power consumption and reduce data rates.
Figure 1: LoRa packet structure
LoRa defines a PHY frame structure, as shown in Fig. 1. The optional header includes the payload length and coding rate and is mandatory for LoRaWAN frames. The optional 16-bit CRC field is required for uplink LoRaWAN frames but is not used for downlink.
The preamble upchirps mark the beginning of a LoRa frame. The number of upchirps is configurable, ranging from 6 to 65,535 (8 for LoRaWAN), while the downchirps are 2.25 symbols long. The network identifier, which is 2 symbols long, distinguishes frames from different networks. Therefore, the minimum length of the LoRa preamble is 10.25 symbols. As a result, using a conventional LoRa chip as a jammer produces a jamming signal that is longer than necessary to achieve the intended jamming objectives, leading to prolonged jammer exposure. Additionally, this approach lacks the flexibility needed to study jamming effects with shorter frames.
In contrast, years of reverse engineering have revealed key details of the LoRa transceiver chain [15, 16]. These efforts have enabled the development of fully functional open-source LoRa transceivers using software-defined radios (SDRs), providing greater flexibility to experiment with LoRa-like modulation without adhering to the conventional frame structure. This method is employed in the setup used in this paper, where jamming LoRa symbols are generated and transmitted as-is, reducing the air-time of the jamming signal and minimizing the risk of exposure, while allowing control over the number of symbols included in the jamming signal.
IV Jamming Analysis Against LoRa protection Mechanisms
Figure 2: LoRa transceiver chain
Figure 3: LoRa diagonal interleaving: (a) before, (b) after; $m=$ SF, $l=4+$ CR
Figure 4: LoRa interleaving example with SF 7 and CR 1
In this section, we discuss the protection mechanisms LoRa provides against transmission errors and examine the impact of jamming on the decoding of received frames with respect to these mechanisms.
The LoRa transceiver chain, as shown in Fig. 2, consists of several components that work together to detect and correct transmission errors. The Hamming encoding/decoding stage provides initial error protection through forward error correction (FEC) codes. These codes are associated with four coding rates (CR): 1, 2, 3, and 4, which correspond to Hamming codes (4,5), (4,6), (4,7), and (4,8), respectively. CR 1 and 2 can only detect errors, while CR 3 and 4 can correct a single-bit error per codeword. This protection is further enhanced by the interleaving/deinterleaving stage. The interleaving process spreads the bits of a codeword across multiple symbols, which, in combination with Hamming encoding, strengthens the frame’s resilience to burst errors.
The interleaving stage operates as a block transformation. It takes $m$ SF codewords, each of length $l=4+$ CR bits, output by the Hamming encoding block, and produces $l$ symbols of $m$ bits. Reverse engineering has revealed that LoRa uses a diagonal interleaving scheme, shown in Fig. 3. The input to the interleaver is an $m\times l$ matrix, where each row contains the bits of a codeword. The output is an $l\times m$ matrix, where the $i$-th row corresponds to the $j$-th column of the input matrix, starting with the last bit and proceeding downwards, followed by an $i$-fold right rotation. The interleaved output symbols contain exactly one bit from each input symbol. An example of this diagonal interleaving for SF 7 and CR 1 is shown in Fig. 4.
For a jamming attack to be effective, the jamming signal must be difficult for the LoRa receiver to mitigate. Several studies have examined the impact of various interference sources on LoRa [17, 18, 19]. These studies show that LoRa transmissions are most affected by interference from another LoRa transmission using the same SF on the same channel, referred to as intra-SF/same-SF interference. Instead of structuring the jamming signal according to the LoRa packet format, we propose using a series of randomly modulated LoRa symbols. This approach allows us to control the length of the jamming signal and study the impact of the number of symbols on the legitimate signal received, as well as determine the minimum number of symbols needed to completely jam the legitimate signal under specific physical parameters.
A reactive jamming scenario is considered, where the jamming signal is assumed to have sufficient power to disrupt the legitimate signal at the receiver.
For the interleaving stage, we assume CR $\in{1,2,3,4}$ and SF $\in{7,8,\dots,12}$. Let $i\in{0,1,\dots,l-1}$ and $j\in{0,1,\dots,m-1}$. $n_{s}$ denotes the number of jamming symbols, where $n_{s}\in[0,CR+4]$.
The first jamming symbol is assumed to be perfectly aligned in time with the first symbol of the block from the legitimate transmitter. The jamming symbol is modulated in the same way as standard LoRa symbols, containing SF bits, which are transmitted as legitimate symbols. The interleaved symbols from the legitimate signal and the jamming symbols are denoted $S^{(L)}$ and $S^{(A)}$, respectively. Specifically, $S^{(L)}{ij}$ and $S^{(A)}{ij}$ represent the $j$-th bit of the $i$-th legitimate and jamming symbols. The whitening stage allows us to assume that the transmitted symbols are random. Similarly, the jamming symbols are generated randomly, following a uniform distribution between 0 and $12^{SF}-$ 1. Thus, both the jamming and legitimate symbol bits can be considered equiprobable: $p(S^{(L)}{ij}=0)=p(S^{(L)}{ij}=1)=0.5$, and the same holds for $S^{(A)}_{ij}$.
The probability that $S^{(L)}{ij}$ is decoded correctly (i.e., not altered by the jammer) is denoted as $p{ij}$, while $\overline{p_{ij}}$ represents the probability that it is jammed (i.e., the decoded value differs from the transmitted value). Hence,
$\begin{split}\overline{p_{ij}}&=P(S^{(L)}_{ij}=0)\times P(S^{(A)}_{ij}=1)\\ &\hskip 8.53581pt+P(S^{(L)}_{ij}=1)\times P(S^{(A)}_{ij}=0)\\ &=0.5\end{split}$ (1)
and therefore,
$p_{ij}=0.5$ (2)
For the remainder of this paper, we define $p=p_{ij}$ and $\overline{p}=\overline{p_{ij}}$, where $P_{suc}$ and $P_{jam}$ represent the probabilities of the frame being successfully decoded and being jammed, respectively. The following analysis does not consider errors caused by noise or interference sources other than the jammer.
IV-A Coding rates 1 & 2
Since CR 1 and 2 do not provide error correction capabilities, a single error in any of the transmitted bits is sufficient to corrupt the entire frame. Therefore, $P_{suc}$ can be determined by considering the number of jamming symbols $n_{s}$: the frame is successfully decoded if none of the $n_{s}$ jamming symbols causes a bit error.
$\begin{split}P_{suc}&=\prod_{k=0}^{n_{s}-1}p^{m}\\ &=p^{m\times n_{s}}\end{split}$ (3)
IV-B Coding rates 3 & 4
We define $C^{(L)}_{j}$ and $C^{(R)}_{j}$ as the $j$-th codeword of the legitimate transmitter before the interleaving stage, and the corresponding codeword at the receiver after the deinterleaving stage, respectively. Since CR 3 and 4 can correct one bit error per codeword, a block is successfully decoded if none of the codewords within that block is corrupted beyond the code’s correction capacity. Therefore, each pair $(C^{(R)}_{j},C^{(L)}_{j})$ has a Hamming distance of at most 1. The probabilities of having 0 or 1 corrupted bit are given by Eq. (4):
$\begin{split}Pr\left(d(C^{(R)}_{j},C^{(L)}_{j})=0\right)&=p^{n_{s}}\mbox{ and,}\\ Pr\left(d(C^{(R)}_{j},C^{(L)}_{j})=1\right)&=\left(n_{s}\atop 1\right)p\cdot\overline{p}\;^{n_{s}-1}\\ &=n_{s}p\cdot\overline{p}\;^{n_{s}-1}\end{split}$ (4)
Therefore, the probability of correctly decoding $C^{(R)}_{j}$ is given by:
$\begin{split}Pr\left(d(C^{(R)}_{j},C^{(L)}_{j})\leq 1\right)&=Pr\left(d(C^{(R)}_{j},C^{(L)}_{j})=0\right)\\ &\hskip 5.69054pt\mbox{ }+Pr\left(d(C^{(R)}_{j},C^{(L)}_{j})=1\right)\\ &=\mbox{ }\overline{p}\;^{n_{s}-1}(\overline{p}+n_{s}\cdot p)\end{split}$ (5)
Finally, the block can be correctly decoded with a probability of $P_{suc}$:
$\begin{split}P_{suc}&=\prod^{m-1}_{j=0}Pr\left(d(C^{(R)}_{j},C^{(L)}_{j})\leq 1\right)\\ &=\overline{p}\;^{m(n_{s}-1)}(\overline{p}+n_{s}\cdot p)^{m}\end{split}$ (6)
IV-C Comparison of coding rates performances against jamming
To compare the performance of CRs 1&2 and 3&4 under jamming, we plot in Fig. 5 the theoretical frame success rate for both CRs, across different jamming signal sizes (in symbols) and the six SF values.
We limit our theoretical analysis to a single interleaved block, so the maximum number of jamming symbols is 5, 6, 7, and 8 for CRs 1, 2, 3, and 4, respectively. While higher numbers could be achieved by extending the analysis with multiples of the single block’s maximum plus a remainder, we restrict the number of jamming symbols to 5 for all CRs to maintain consistency with the single block analysis.
Under the same SF and number of jamming symbols, the advantage of the error correction capability of CR 3&4 is evident: the frame success rate with CR 3&4 is at least 100 times higher than that of CR 1&2. Furthermore, for CR 3&4 with 1 jamming symbol, the frame success rate is 1 for all SFs, which is expected. Under the assumption that the jamming and legitimate symbols are perfectly aligned, one jamming symbol can at most corrupt 1 bit of each codeword in the block, and any corrupted codeword can be corrected through the error correction capability. Additionally, for the same CR and number of jamming symbols, higher SFs result in lower frame success rates—with a two-fold decrease for each SF increase in CR 1&2. This is expected because higher SF symbols contain more bits, increasing the probability of bit corruption and decreasing the likelihood of successful frame decoding.
The most striking observation from this figure is the effectiveness of the jamming: with only 5 jamming symbols, the frame success rate for CR 3&4 drops to $10^{-5}$ for SF7 and $10^{-9}$ for SF12. Nearly identical results are achieved for CR 1&2 with just 2 jamming symbols. Compared to conventional LoRa chips, which use a 10.25-symbol preamble and no payload, this represents a reduction in jamming symbol count by a factor of two for CR 3&4 and five for CR 1&2, thus reducing jammer exposure time. This demonstrates how harmful this jamming can be to LoRa with very little transmission time.
Figure 5: Theoretical frame success rate under jamming: (a) CR 1&2, (b) CR 3&4
V Experimental assessment
The GNU Radio implementation developed for the experiment is hosted on GitHub at [20] for research purposes only, with the important disclaimer that any form of jamming is illegal. The reactive jammer is divided into three complementary blocks:
(i) FrameDetect block: This block is responsible for detecting the beginning of a LoRa transmission, using physical layer parameters such as the central frequency, bandwidth, and SF. Upon detecting a frame, FrameDetect attempts to recover its header and checks the header’s CRC for validity. If the header is valid, it sends a message containing the SF and CR to the trigger block.
(ii) Trigger block: This block determines the number of jamming symbols to generate for each CR value. When it receives a message from the FrameDetect block, it randomly generates symbol values according to the prior configuration and sends them to the jammer block.
(iii) Jammer block: This block modulates the received symbol values according to its physical layer configuration (SF, bandwidth, sampling rate) and transmits the modulated symbols to the next block, which could be an RF transmitter.
To address the synchronization issue identified in [14], a symbol repetition method is used. If the configured number of jamming symbols exceeds one, the trigger block generates the first symbol value randomly and repeats it for the required number of times.
The experimental setup is shown in Fig. 7. The transmitting end node consists of a USRP-2920 connected to a Raspberry Pi 4 running the LoRa implementation from [16]. The jammer is made up of a USRP-2920 connected to a 3 GHz, 6-core laptop running our jammer implementation, and the receiver is a TTGO-LoRa32.
The end node transmits at 10 dBm, while the jammer is configured with a transmission power of 13 dBm. The experiment targets SF values of 7 and 10, and 1000 frames are transmitted for each combination of CR and the number of jamming symbols. The resulting frame success rates are displayed in Figs. 8a and 8b, along with their theoretical counterparts. The experimental frame success rate’s lower bound is 0 frames received out of the 1000, but for plotting purposes, we set this lower bound to $10^{-5}$ on the logarithmic scale.
For both SF values, the plots corresponding to CR 1 and 2 are very close to each other, as are those for CR 3 and 4. This confirms that CR 1 and 2 deliver the same performance regarding error correction, as do CR 3 and 4.
Additionally, the experimental curves closely follow the theoretical ones, with a slight deviation for CR 3&4 and 1 jamming symbol. This difference arises from the initial assumption that the jamming symbol is perfectly aligned with the legitimate symbol, which is not the case in practice. Instead, the jamming symbol is split and affects two successive legitimate symbols, potentially corrupting more than one bit per codeword.
Figure 6: LoRaJam GNURadio implementation flowgraph
Figure 7: Experiment bench equipment : (a) Top view disposal, (b) TTGO-LoRa32 receiver, (c) USRP 2920 + Raspberry Pi 4 transmitter, (d) USRP 2920 + Laptop as Jammer
Figure 8: Experimental Frame success rate under jamming: (a) SF7, (b) SF10
TABLE I: Signal to Jamming Ratio (in dB)
CR1 CR2 CR3 CR4
SF7 18.42 19.17 16.79 17.35
SF10 13.18 13.33 12.60 13.09
TABLE II: Jamming symbol gain (in dB)
CR1&2 CR3&4
SF7 9.6 6.59
SF10 9.6 7.84
Finally, Tables 1 and 2 demonstrate the effectiveness of our jamming approach. Table 1 presents the signal-to-jamming ratio (SJR) for each combination of spreading factor (SF) and coding rate (CR). It is calculated as the ratio of the number of legitimate symbols, weighted by transmission power, to the number of jamming symbols, weighted by jamming power. The legitimate symbols correspond to physical payloads of 180 bytes for SF7 and 60 bytes for SF10, both of which are below the maximum allowable payload size. Table 2, on the other hand, shows the gain in jamming symbols achieved by our method compared to a conventional LoRa transceiver, which transmits 18.25 symbols without a payload. The computed gains would be even greater if the payload were included for the conventional case. This once again highlights how susceptible the standard LoRa modulation is when faced with a well-designed jammer.
VI Conclusion
In this paper, we explored reactive jamming attacks against LoRaWAN, focusing on minimizing jammer exposure time. We identified and analyzed the key components of the LoRa transceiver chain that mitigate jamming attacks. Our analysis demonstrates how the frame success rate is impacted by the number of jamming symbols transmitted. It also shows that transmitting fewer than one interleaving block is enough to significantly reduce the frame success rate, provided the jamming power is sufficient for the jammer’s symbol to be misinterpreted as the legitimate symbol by the receiver. Additionally, we presented a GNU Radio implementation of the jammer to validate our analysis. The deterministic and known preamble pattern of LoRa makes reactive jamming feasible and straightforward. Future work will focus on developing a stealthy preamble structure for LoRa as a countermeasure to reactive jamming.
Acknowledgment
This work was sponsored by the Junior Faculty Development program under the UM6P-EPFL Excellence in Africa Initiative.